OpenStudy (anonymous):

Do not allow to post question if it doesn't meet certain conditions. For example every question should be 30 (or any other reasonable number) characters long. It would get rid of questions such as: "please help", "see atachment" and etc.

OpenStudy (anonymous):

maybe but I think the limit should be a bit shorter, say, 15 words.

OpenStudy (anonymous):

15 words is much more than 30 characters :P usually

OpenStudy (anonymous):

oh, I thought you meant 30 words. yes, 30 characters would be correct.

OpenStudy (aroub):

That's a not good idea... ( sorry )

OpenStudy (aroub):

like for example ' see attachment' it's important what else should you write?

OpenStudy (anonymous):

You can describe what's in attachment, attachments should be only geometric figures or graphs. But other people post whole problems since they are too lazy to type and you don't even know about what is that problem so you need to open question, open attachment and see that it's calculus which you might haven't learnt.

OpenStudy (aroub):

Okay, then let them write see attachment calculus or something. I hate when some websites have a limit for words! maybe I want to write less or more...

OpenStudy (anonymous):

word calculus is not descriptive, it's ALWAYS better to describe problem as much as possible and you SHOULD explain what you have done, not just say: Calculus, see attachment, give answer, kthxbye

OpenStudy (aroub):

Haha! yes I think you're right but what should they write? describe the question? like: ( not calculus lol) this question is algebra1 it contains three parts you have to find x for the first part and the second part...blah blah,but it's just better if you wrote 'Algebra 1( or 2..), see attachment. Please help' then who took algebra 1 can help NOT DESCRIBE THE QUESTION! right? I really don't want to argue on this.. maybe you're like 80% right!

OpenStudy (anonymous):

They can write, i was trying to find x but by multiplying both sides by 5 and I have done it but answer is still wrong look x/5=10 (multiply by 5) x/5 * 5=10 x = 10 wheres my mistake? and then I come and say oh dear you forgot to multiply other side it should have been x/5 * 5 = 10 * 5 x = 50 TADAM!

OpenStudy (anonymous):

I disagree

OpenStudy (preetha):

I think there should be a guidelines on both questions and answers. But when would we put that to new users?

OpenStudy (anonymous):

I don't know you can write about it in Terms and Conditions (which is not read anyway :P) but the point is that system would automatically spawn message that your question is too short and it wouldn't allow to post question until you add more characters. So you won't need any human power to control this. And I am pretty sure this can be easily done, you can ask your programming colleagues :P And others disagree because they like to do it themselves..?

OpenStudy (anonymous):

Placing controls on question length not a good idea. However, providing guidelines (offline and infrequent inline reminders of such guidelines) are more helpful. Placing character and word count limits on questions is just not good. For example: "Solve x^2+8x+4" is enough of a problem statement. But, may not meet the "character / word" threshold.

OpenStudy (anonymous):

You didn't read what I say, person SHOULD show what he has done "solve x^2+8x+4" doesn't show anything only that he's so lazy that need only answer and can't even google quadratic equation solver or go to wolframalpha or whatever

OpenStudy (aroub):

Aooh! I thought to DESCRIBE the question lol okay then I agree with you :) but.. idk if that will work 'to limit the number of characters'

OpenStudy (anonymous):

of course you can get around of it but it would be better

OpenStudy (underhill):

That wouldn't be a good idea - it would place emphasis where emphasis doesn't belong. We really don't care how long the question is - just how good it is. This is a study group, so its true that people shouldn't be looking for a cut-and-dried answer to a word-for-word problem. (I've seen one person ask for an answer to a test-question while taking the test!) But that's a problem that should not be solved by limiting the characters in a question.